OpenStudy (anonymous):

Write the equation of the line that passes through the points (–1, –4) and (2, 5). Write the equation of the line that passes through the points (–1, –4) and (2, 5). @Mathematics

OpenStudy (turingtest):

\[m={y_2-y_1 \over x_2-x_1}={5-(-4) \over 2-(-1)}=3\]\[y-y_1=m(x-x_1)\]\[y-5=3(x-2)\]\[y=3x-1\]